Multiply.
(sqrt11 - sqrt5)(sqrt11 + sqrt5)
Thanks

Answer by fractalier(6550)

Like any (a+b)(a-b) = a^2 - b^2 we get
(sqrt11 - sqrt5)(sqrt11 + sqrt5) =
11 - 5 =
6